OpenStudy (anonymous):

f(n)=sin4x+cos4x local max and min?

OpenStudy (anonymous):

work with \[\sqrt{2}\sin(x+\frac{\pi}{4})\]

OpenStudy (anonymous):

oops it is 4x, ok \[\sqrt{2}\sin(4x+\frac{\pi}{4})\]

OpenStudy (anonymous):

so max is \[\sqrt{2}\] and min is \[-\sqrt{2}\]

OpenStudy (anonymous):

The answer appears to be local min pi/4, 3pi/4, 5pi/4, 7pi/4 and local max; pi/2 and 3pi/2

OpenStudy (anonymous):

i will respectfully disagree

OpenStudy (anonymous):

first of all max in an output, not an input

OpenStudy (anonymous):

shoot, im an idiot. I meant \[f(n)=\sin^4x+\cos^4x \]

OpenStudy (anonymous):

oooooooooooooooooooh

OpenStudy (anonymous):

then i still respectfully disagree as for the reason above, max is an output not an input it will have a max when \(x=\frac{\pi}{4}\) and at that point you will get 1, so the max is one

OpenStudy (anonymous):

sorry what you said, max at \(x=\frac{\pi}{2}\) but the max is 1

OpenStudy (anonymous):

but my calculus textbook! it says otherwise. Can you show me the method?